OpenStudy (anonymous):

How many different 6-digit numbers can be formed using the digits 1,2,3,4,5,6 if repetitions are allowed?

jimthompson5910 (jim_thompson5910):

You have 6 choices for each slot, and you have 6 slots, so you have 6*6*6*6*6*6 = 6^6 = 46,656 different 6-digit numbers

OpenStudy (anonymous):

thank you so much

jimthompson5910 (jim_thompson5910):

np